# On the Conditions for a Quantum Violent Relaxation

### 22 Dec 2023 v1

@@ Abstract
Violent relaxation is a characteristic feature of classical long-range systems out of equilibrium [23, 1]. This phenomenon has been observed in fully-connected models, including the paradigmatic Hamiltonian-Mean-Field (HMF) model [24, 1], and it is believed to be a key ingredient to explain the formation of large-scale astrophysical structures [25, 26, 27, 28, 10, chavanis1998degenerate, chavanis2005coarse, chavanis2006coarse]. The HMF dynamics, in the thermodynamic limit, becomes mean-field, thus preventing the loss of memory and the onset of the thermalization. In spite of this, on a finite (size-independent) time-scale, the system is usually observed to relax toward a so-called quasi-stationary state, in which large-scale observables take a constant (non-thermal) value. While the general mechanism behind the phenomenon is generally understood [2, 26, 28, 29, chavanis2006lynden, chavanis2006coarse], no general framework, comparable to the H-theorem in the usual kinetic theory [30], has emerged yet.
In the quantum regime, fully connected Hamiltonians, akin to the HMF, can be engineered in cavity experiments [17, 31] or Rydberg gases [32, 33] or seen as a large-dimension approximation of short-range models [34, 35]. These models display a rich out-of-equilibrium phenomenology, including dynamical phase transitions [36, 37, 38, 39], anomalous entanglement scaling [40, 41, 42], Floquet time-crystals [43, 44, 45], which has proven to be robust against external noise and competing short-range perturbations [15, 46, 16, 47]. As the mean-field description becomes exact [pearce1975anisotropic, pearce1978high, tindemans1974exact, kastner2010nonequivalence], according to the general wisdom, the dynamics of fully-connected systems can be reduced to a small number of classical macroscopic degrees of freedom in the thermodynamic limit [37, 38], resulting in long-lived periodic oscillations of the order parameter (a picture easily generalized to most strong-long range systems [16]). This is the case, e.g., of the paradigmatic Lipkin-Meshkov-Glick (LMG) model [48, 39].
@@
where the sum runs all over the permutations. In the thermodynamics limit, the mean-field description of the system becomes exact [pearce1975anisotropic, pearce1978high, tindemans1974exact, kastner2010nonequivalence, 37, 38]. This means that, as long as we are interested in single-site operators, the dynamics of the system can be described (see [55] for a physical argument) in terms of a single-site effective density operator $\rho(t)$, evolving through the Hamiltonian

### 14 Jan 2024 v2

@@ Abstract
Violent relaxation is a characteristic feature of classical long-range systems out of equilibrium [23, 1]. This phenomenon has been observed in fully-connected models, including the paradigmatic Hamiltonian-Mean-Field (HMF) model [24, 1], and it is believed to be a key ingredient to explain the formation of large-scale astrophysical structures [25, 26, 27, 28, 10, 29, 30, 31]. The HMF dynamics, in the thermodynamic limit, becomes mean-field, thus preventing the loss of memory and the onset of the thermalization. In spite of this, on a finite (size-independent) time-scale, the system is usually observed to relax toward a so-called quasi-stationary state, in which large-scale observables take a constant (non-thermal) value. While the general mechanism behind the phenomenon is generally understood [2, 26, 28, 32, 33, 31], no general framework, comparable to the H-theorem in the usual kinetic theory [34], has emerged yet.
In the quantum regime, fully connected Hamiltonians, akin to the HMF, can be engineered in cavity experiments [17, 35] or Rydberg gases [36, 37] or seen as a large-dimension approximation of short-range models [38, 39]. These models display a rich out-of-equilibrium phenomenology, including dynamical phase transitions [40, 41, 42, 43], anomalous entanglement scaling [44, 45, 46], Floquet time-crystals [47, 48, 49], which has proven to be robust against external noise and competing short-range perturbations [15, 50, 16, 51]. As the mean-field description becomes exact [52, 53, 54, 55], according to the general wisdom, the dynamics of fully-connected systems can be reduced to a small number of classical macroscopic degrees of freedom in the thermodynamic limit [41, 42], resulting in long-lived periodic oscillations of the order parameter (a picture easily generalized to most strong-long range systems [16]). This is the case, e.g., of the paradigmatic Lipkin-Meshkov-Glick (LMG) model [56, 43].
@@
where the sum runs all over the permutations. In the thermodynamics limit, the mean-field description of the system becomes exact [52, 53, 54, 55, 41, 42]. This means that, as long as we are interested in single-site operators, the dynamics of the system can be described (see [63] for a physical argument) in terms of a single-site effective density operator $\rho(t)$, evolving through the Hamiltonian